OpenStudy (anonymous):

The area of a rectangle is x^2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle. A. x - 5 B. x^2 + 15 C. x + 5 D. (x - 2)(x + 3)

OpenStudy (anonymous):

Area = l x w x² - 2x - 15 = (x+3)w (x²-2x-15)/(x+3) = w w= ??